OpenStudy (anonymous):

What is the equation of the line that passes through the points (−1, 3) and (1, 11)?

Nnesha (nnesha):

do you know the formula to find slope ?

OpenStudy (anonymous):

y=mx+b

Nnesha (nnesha):

that's the equation we need it later but first we have to find slope \[\huge\rm \frac{ y_2 -y_1 }{ x_2 -x_1 }\] and plug in order pair where x and y are \[\huge\rm (x_1 ,y_1) (x_2 ,y_2) \]

OpenStudy (anonymous):

so it would be 11-3 over 1-(-1)

Nnesha (nnesha):

yes yes right !! :=) solve

OpenStudy (anonymous):

8 over 2

OpenStudy (anonymous):

then it would be 4

Nnesha (nnesha):

yes right now use that equation plug in m balue m=slope and pick one of order pair replace xand y by their values solve for b

OpenStudy (anonymous):

ok thank u

Nnesha (nnesha):

let me know what you get!

OpenStudy (anonymous):

i dont understand what u said^

Nnesha (nnesha):

cool and thanks for telling me :=) so \[\huge\rm y = mx+b\] m=4 now you can pick one of order pair where x and y values are (x,y) and then substitute into the equation let's say (1,11) so replace x by 1 y by 11 and m by 4 solve for b

OpenStudy (anonymous):

7?

Nnesha (nnesha):

\[\huge\rm \color{red}y=\color{green}{m}\color{blue}{x}+b\] \[(\color{blue}{1},\color{Red}{11})\]

OpenStudy (anonymous):

is the answer 7

Nnesha (nnesha):

yes b =7 \[\huge\rm y=\color{green}{m}{x}+\color{ReD}{b}\] m=4 b=7 plug in &&& done!

OpenStudy (anonymous):

omg thanks

Nnesha (nnesha):

my pleasure